OpenStudy (anonymous):

What is the VALUE OF:

OpenStudy (anonymous):

\[\huge \left( \frac { 81 }{ 16 } \right) ^{ \frac { -3 }{ 4 } x }\]

OpenStudy (anonymous):

any one would like to have this

OpenStudy (anonymous):

\[\And \ also \ this \\ \huge \left( \frac { 25 }{ 9 } \right) ^{ \frac { -3 }{ 2 } }\]

OpenStudy (anonymous):

\[\huge ( \frac{3}{2})^{4 \times \frac{-3x}{4}} \implies (\frac{3}{2})^{-3x}\]

OpenStudy (anonymous):

Which one have you done @waterineyes

OpenStudy (anonymous):

\[\And \ also \ this \\ \huge \left( \frac { 25 }{ 9 } \right) ^{ \frac { -3 }{ 2 } } \implies (\frac{5}{3})^{2 \times \frac{-3}{2}} \implies (\frac{5}{3})^{-3}\]

OpenStudy (anonymous):

Good attempt @waterineyes

OpenStudy (anonymous):

For first one: \[\large (\frac{3}{2})^{-3x} \implies (\frac{2}{3})^{3x} \implies (\frac{8}{27})^{x}\]

OpenStudy (anonymous):

Okay! I knew the answer but.... still my method is far different from yours!

OpenStudy (anonymous):

For second one: \[\large (\frac{5}{3})^{-3} \implies (\frac{3}{5})^3 \implies \frac{27}{125}\]